A truck has a maximum acceleration of 2.0 m/s^2 when its total mass is 1800 kg. What is the max acceleration of the truck after picking up an additional load of 400 kg
Answer #1

Force exerted by engine is same as F = ma

m1a1 =m2a2

1800*2 = (1800+400)*a2

a2 = 1800*2/2200

= 1.6363 m/s^2